OpenStudy (anonymous):

Find the solution of this system equations. -2x+8y=14 and -2x-2y=-26 Help? I don't get this at all.

OpenStudy (anonymous):

First put in to slope intercept form

OpenStudy (anonymous):

then divide by y

OpenStudy (anonymous):

-2x + 8y = 14 -2x - 2y = -26 --->(-1) -------------------- -2x + 8y = 14 2x + 2y = 26 (result of multiplying by -1) ---------------------add 10y = 40 y = 40/10 y = 4 now sub in 4 for y in either of the equations to find x -2x + 8y = 14 -2x + 8(4) = 14 -2x + 32 = 14 -2x = 14 - 32 -2x = -18 x = -18/-2 x = 9 check.. -2x + 8y = 14 -2(9) + 8(4) = 14 -18 + 32 = 14 14 = 14 (correct) x = 9 and y = 4 or (9,4)

OpenStudy (anonymous):

Thank you so much! Could you help me with others? @kelliegirl33